# Video Soundtrack Generation by Aligning Emotions and Temporal Boundaries

### v1

@@ Abstract
We introduce EMSYNC, a video-based symbolic music generation model that aligns music with a video’s emotional content and temporal boundaries. It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues. We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate and align musical chords with scene cuts. Unlike existing models, our approach retains event-based encoding, ensuring fine-grained timing control and expressive musical nuances. We also propose a mapping scheme to bridge the video emotion classifier, which produces discrete emotion categories, with the emotion-conditioned MIDI generator, which operates on continuous-valued valence-arousal inputs. In subjective listening tests, EMSYNC outperforms state-of-the-art models across all subjective metrics, for music theory-aware participants as well as the general listeners.
@@ 1 Introduction
The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software Falkowski-Gilski and Uhl (2020). A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement Buhler (2018). However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube. Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content. Automatic video-based music generation offers a promising solution to this problem Di et al. (2021); Kang et al. (2024); Zhuo et al. (2023).
@@
Existing approaches that generate music as audio waveforms lack editability Lam et al. (2024); Copet et al. (2024). This method compresses all stages of music production—composition, performance, editing, mixing, and mastering—into a single process, limiting creative control for professionals. By contrast, generating music in a symbolic format like MIDI (Musical Instrument Digital Interface) offers greater flexibility. MIDI functions as a digital music score, encoding instrument names, note pitches, durations, and velocities. Professionals can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
@@
This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs). We use “MIDI” and “music” interchangeably, as MIDI is our exclusive format. A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets. Existing datasets are either domain-specific, such as pianist hand videos Koepke et al. (2020), or contain limited samples, around 1k Di et al. (2021); Zhuo et al. (2023). To address this, we develop a model that generates music for any video type, leveraging the Lakh MIDI Dataset—the largest available MIDI dataset with 176,581 samples—ensuring diverse and high-quality outputs Raffel (2016).
@@
Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs. We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries. While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model Russell (1980). Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near those locations that are rhythmically and harmonically compatible with the rest of the generated music. We also use a pretrained video emotion classifier to estimate discrete emotion probabilities Sulun et al. (2024a), map them to valence-arousal values, and condition our music generator accordingly.
@@
Temporal conditioning in MIDI-generating models presents unique challenges. While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate Liu et al. (2022); Arnab et al. (2021). In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated Oore et al. (2020). In event-based MIDI encoding, two primary token types exist: “note” and “time shift”. Note tokens represent pitch, while time shift tokens advance the time axis, capturing both note durations and silences. Each time shift token specifies a time increment, forming a one-dimensional sequence where the position in the sequence does not directly correspond to the position in time. The key advantage of event-based encoding is the absence of a fixed time grid, allowing for subtle, expressive timing variations that reflect human musicianship. Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids Di et al. (2021); Kang et al. (2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
We also tackle the challenge of using emotions as an intermediary to link music generation to video. Sulun et. al Sulun et al. (2022) previously assigned valence-arousal emotion labels to samples in the Lakh MIDI dataset Raffel (2016), creating a training dataset two orders of magnitude larger than existing emotion-labeled MIDI datasets Panda et al. (2013); Ferreira and Whitehead (2019); Hung et al. (2021). However, the largest emotion-labeled video datasets primarily provide categorical labels Xu et al. (2016); Sun et al. (2020). To bridge this gap, we employ the VEMOCLAP model (Video EMOtion Classifier using Pretrained features) Sulun et al. (2024a), trained on the Ekman-$6$ dataset Xu et al. (2016). Since VEMOCLAP outputs probabilities of categorical emotions, we map these to valence-arousal values using prior user study results on emotion Russell and Mehrabian (1977).
We compare our method to state-of-the-art video-based MIDI generators through subjective listening tests, where it outperforms all competing methods across all metrics for both music theory-aware participants and general listeners while also being the fastest. Our key contributions include the following:
@@
  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective metrics and participant demographics.
@@
  • We leverage insights from psychological studies to map discrete emotion categories to continuous valence-arousal values, enabling an emotional connection between the input video and the generated music.
  • We design our model architecture to modify only the inputs to the transformer body, allowing future users to fine-tune it for other tasks and data structures seamlessly.
@@ 2 Related work
In this section, we introduce the video emotion classifier and the emotion-based MIDI generator used in our work. We then review existing video-based MIDI generation methods and discuss their evaluation approaches.
2.1 Video emotion classification
The VEMOCLAP model exploits publicly available pretrained models for a multimodal video analysis and uses the resulting pretrained features to classify the emotions of arbitrary videos Sulun et al. (2024a, b). It employs pretrained models for automatic speech recognition (ASR), optical character recognition (OCR), facial expression classification, audio classification, and image understanding. Textual features obtained from ASR and OCR are further processed using a text sentiment classifier. Cross-attention layers Vaswani et al. (2017) integrate and process these multimodal features, and a linear layer produces the final emotion probabilities. The model is trained on the Ekman-6 dataset Xu et al. (2016), consisting of 1637 videos labeled with the six basic emotions derived from the original work of Ekman: anger, disgust, fear, joy, sadness, and surprise Ekman (1971).
2.2 Emotion-based MIDI generation
Sulun et al. Sulun et al. (2022) collected valence and arousal values for songs in the Lakh Pianoroll Dataset Dong et al. (2018) and trained a conditional transformer using the resulting labeled dataset. To integrate continuous-valued valence and arousal features with discrete musical note tokens, they project valence and arousal into a vector space using separate linear layers. Musical input tokens are projected into vectors of the same dimensionality using an embedding layer. These vectors are then concatenated with the projected valence and arousal vectors along the sequence dimension and fed into the transformer body with relative global attention Huang et al. (2018).
2.3 Video-based MIDI generation
@@
While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances. The Foley Music model Gan et al. (2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network Kipf and Welling (2017) and a Transformer Vaswani et al. (2017). Similarly, Koepke et al. Koepke et al. (2020) and Su et al. Su et al. (2020) utilize use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards. Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, though these datasets typically contain fewer than 1k samples Zhao et al. (2018); Li et al. (2018); Koepke et al. (2020); Su et al. (2020). The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music Su et al. (2021).
@@
Some studies explore the more general task of generating symbolic music for arbitrary videos. The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing Di et al. (2021). CMT employs the Lakh Pianoroll Dataset Dong et al. (2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration Hsiao et al. (2021). While this representation reduces sequence length, it significantly increases input dimensionality. Consequently, CMT uses a low temporal resolution of four ticks per beat, translating to a time resolution of 125 milliseconds for a song at 120 beats per minute. Furthermore, note events align precisely with beat subdivisions, whereas human musicians introduce expressive timing deviations. In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8-millisecond resolution. This resolution is sufficiently fine to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions Oore et al. (2020). Additionally, unlike CMT, our model integrates high-level emotional conditioning alongside low-level temporal features.
@@
In a follow-up to CMT, Zhuo et al. introduced the Symbolic Music Videos dataset, containing video-MIDI pairs Zhuo et al. (2023). They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples Hawthorne et al. (2018). Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features Radford et al. (2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
@@
The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning but differs in application Kang et al. (2024). The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key. Their encoder-decoder transformer takes low- and high-level video features, along with user-provided primer chord and key, to generate chord sequences. These sequences are then arpeggiated using fixed patterns to create the final MIDI output. However, this method’s reliance on a fixed time grid omits the subtle expressive timing found in human performances. Additionally, its use of fixed arpeggiation patterns limits musical diversity, and requiring user-defined chord and key inputs restricts accessibility for non-musicians.
2.3.1 Evaluation methods
Evaluation of video-based music generators is non-trivial. For objective evaluation, most works measure the difference between generated and ground-truth samples using metrics such as Pitch Class Histogram Entropy, Grooving Pattern Similarity, and Structureness Indicator Wu and Yang (2020). When the dataset contains paired video and MIDI samples, this method is straightforward Gan et al. (2020); Zhuo et al. (2023). For unpaired video and MIDI samples, objectively evaluating a model that generates MIDI from arbitrary videos is not feasible.
The authors of the Controllable Music Transformer, using unpaired datasets, evaluate output MIDIs generated unconditionally, without input video Di et al. (2021). They empirically show that unconditionally generated MIDI samples exhibit metrics closer to the unpaired MIDI dataset compared to video-conditioned output samples. This occurs because constraining the model with video structure causes deviations from intrinsic MIDI structures, and the best way to match a specific MIDI dataset’s metrics is to train exclusively on that dataset. Subjective evaluation through user studies remains the most common, and often the only, viable method for assessing video-based MIDI generators Di et al. (2021); Kang et al. (2024); Zhuo et al. (2023); Gan et al. (2020); Su et al. (2021).
@@ 3.1 Conditional music generator
Our music generator is conditioned on emotions as valence-arousal values and temporal boundaries. Figure 2 illustrates the model, with the emotion-conditioning mechanism in the upper part and the boundary-conditioning mechanism in the lower part.
@@
3.1.1 Training dataset and preprocessing
We train our music generator on the Lakh Pianoroll Dataset (LPD) Dong et al. (2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset Raffel (2016). We tokenize the pianorolls using an event-based symbolic music representation Oore et al. (2020). Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end. These tokens also encode pitch and instrument information. For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60. Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings Dong et al. (2018). However, our method is adaptable to datasets with different instrument groupings.
TIMESHIFT tokens are used to move along the time axis, representing both note durations and the silences between them. Each token specifies a time increment in milliseconds. For example, an 800-millisecond shift is encoded as TIMESHIFT_800. We use a temporal resolution of 8 milliseconds with a maximum shift of 1000 milliseconds. Longer durations are represented using multiple consecutive TIMESHIFT tokens. We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches. We use the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions Bharucha and Krumhansl (1983). However, this is a design choice, and our method can accommodate any selection of temporal boundaries.
For emotion-conditioned generation, we use the model architecture and valence-arousal labeled dataset of Sulun et al., detailed in Section 2.2, which includes 23,661 samples Sulun et al. (2022). To utilize the remaining unlabeled samples, we further modify the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation. If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
@@ 3.1.1 Training dataset and preprocessing
3.1.2 Temporal boundary matching
In addition to matching emotions, we also temporally synchronize the input video and output music by aligning their temporal boundaries. We define temporal boundaries as scene cut locations in video and chord locations in music. Since we employ a hybrid approach, we first train a music generator capable of incorporating boundary locations as input and producing music with chords near these boundaries. To achieve this, we label the chords in the Lakh Pianoroll Dataset. During training, we consider only guitar and piano chords with at least three simultaneous notes that last for a minimum of two beats. These chords are labeled by inserting a CHORD token before the first ON token of each chord. To allow the model to generate chords not only at video scene cut locations but also independently where musically appropriate, we randomly remove 20% of CHORD tokens during training. An example of a labeled chord is shown in the top row of Figure 3. In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
@@ 3.1.2 Temporal boundary matching
Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following Bharucha and Krumhansl (1983). During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt. This occurs because the model, having no prior knowledge of the upcoming chord, may generate preceding notes that do not align naturally with it. To address this, we develop a method that enables the model to “anticipate” upcoming chords and generate preceding notes and time shifts accordingly. Additionally, we train the model to generate the CHORD token itself, ensuring rhythmic consistency in the generated music. To achieve this, we define boundary offsets for each input token, representing the remaining time until the next boundary. These offsets are capped at a maximum value and, since our music generator is autoregressive, they are computed based on future chords rather than past ones. Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music. Sample offsets are illustrated in the bottom row of Figure 3.
We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings Wang et al. (2020) along the feature dimension, as shown in the lower part of Figure 2. This architectural choice is motivated by several key factors. First, similar to the emotion-conditioning mechanism, we inject boundary offset encodings at the input level rather than within the transformer body, ensuring that the core model remains unchanged. This allows the model to process inputs with or without boundary offsets, enabling seamless fine-tuning by repeating the learned positional encodings along the feature dimension when boundary offsets are absent. Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two. Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them Haviv et al. (2022); Kazemnejad et al. (2024). Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings. The resulting vector sequence consists of positional encodings augmented with boundary offset encodings. Following the standard transformer model, we add this sequence to the token embeddings Vaswani et al. (2017) before passing it to the transformer body with relative global attention Huang et al. (2018). The transformer’s input is represented as:
where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the input boundary offsets, and $\mathbf{W_{pe}}$ is the learned positional encoding. The feed-forward networks $\text{FFN}_{v}$, $\text{FFN}_{a}$, and $\text{FFN}_{b}$ process the valence, arousal, and boundary offset inputs, respectively. The operations $\text{Concat}_{s}$ and $\text{Concat}_{f}$ denote concatenation along the sequence and feature dimensions, respectively. Our music generator architecture is illustrated in Figure 2.
@@
Algorithm 1 outlines the process of computing boundary offsets during inference, i.e., music generation. The model generates music autoregressively, producing one token per forward pass. By tracking the generated TIMESHIFT tokens, we maintain a time cursor and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token. If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary. The boundary with a difference smaller than the predefined sensitivity threshold of 1 second is considered successfully generated. We then remove these boundaries from future offset calculations by replacing them with infinity. Next, we compute the boundary offset for the generated token, regardless of its type. This offset is determined as the distance to the next closest boundary, capped by a maximum offset value. For simplicity, Algorithm 1 is presented for a single sample, but in practice, this operation is performed in minibatches. During training, we preprocess the entire input sequence at once by constructing a time grid instead of a time cursor, allowing us to calculate boundary offsets for all tokens simultaneously. For clarity, the initial list of generated tokens is shown as empty; however, in practice, we begin with a START token.
@@
     $\boldsymbol{\delta}$.append$(\max(\min(\mathbf{b}-c),\delta_{max}))$
     $\mathbf{t}$.append$((t_{t},t_{v}))$

### v2

@@ Abstract
We introduce EMSYNC, a video-based symbolic music generation model that aligns music with a video’s emotional content and temporal boundaries. It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues. We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate upcoming video scene cuts and align generated musical chords with them. Unlike existing models, our approach retains event-based encoding, ensuring fine-grained timing control and expressive musical nuances. We also propose a mapping scheme to bridge the video emotion classifier, which produces discrete emotion categories, with the emotion-conditioned MIDI generator, which operates on continuous-valued valence-arousal inputs. Our method outperforms state-of-the-art models in objective and subjective evaluations across different video datasets, demonstrating its effectiveness in generating music aligned to video both emotionally and temporally.
@@ 1 Introduction
The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software (Falkowski-Gilski and Uhl 2020). A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement (Buhler 2018). However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube. Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content. Automatic video-based music generation offers a promising solution to this problem (Di et al. 2021; Zhuo et al. 2023; Kang, Poria, and Herremans 2024).
@@
Existing approaches that generate music as audio waveforms lack editability (Lam et al. 2024; Copet et al. 2024). These methods compress all stages of music production, including composition, performance, editing, mixing, and mastering, into a single process, which limits creative control for musicians and producers. By contrast, generating music in a symbolic format like MIDI (Musical Instrument Digital Interface) offers greater flexibility. MIDI functions as a digital music score, encoding instrument names, note pitches, durations, and velocities. Musicians and producers can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
@@
This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs). We use “MIDI” and “music” interchangeably, as MIDI is our exclusive format. A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets. Existing datasets are either domain-specific, such as pianist hand videos (Koepke et al. 2020), or contain limited samples, around 1k (Di et al. 2021; Zhuo et al. 2023). To address this, we develop a model that generates music for any video type by leveraging the Lakh MIDI Dataset (Raffel 2016), the largest available collection with 176,581 samples, to ensure diverse and high-quality outputs.
@@
Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs. Drawing from musicology literature, we identify two essential aspects for effective video-based music generation: emotional alignment (Deutsch 2007) and temporal synchronization (Prendergast 1992). We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries. While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model (Russell 1980). Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near those locations that are rhythmically and harmonically compatible with the rest of the generated music. We also use a pretrained video emotion classifier to estimate discrete emotion probabilities, map them to valence-arousal values, and condition our music generator accordingly.
@@
Temporal conditioning in MIDI-generating models presents unique challenges. While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate (Arnab et al. 2021; Liu et al. 2022). In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated (Oore et al. 2020). In event-based MIDI encoding, two primary token types exist: “note” and “time shift”. Note tokens represent pitch, while time shift tokens advance the time axis, capturing both note durations and silences. Each time shift token specifies a time increment, forming a one-dimensional sequence where the position in the sequence does not directly correspond to the position in time. The key advantage of event-based encoding is the absence of a fixed time grid, allowing for subtle, expressive timing variations that reflect human musicianship. Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids (Di et al. 2021; Kang, Poria, and Herremans 2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
The selection of temporal features for video-music synchronization is also non-trivial. Existing work matches video motion speed and motion saliency with musical note density, producing a dense representation of temporal dynamics that leads to continuous conditioning and frequent changes in note density (Di et al. 2021). Our inspection of the output samples from this approach identified two key limitations. First, the continuously varying note density can disrupt rhythmic consistency, making the underlying beat unstable and harder to follow. Second, because the mapping is dense, synchronization between music and video is difficult to perceive unless there are strong contrasts in motion speed and note density across consecutive sections. Motivated by works in musicology and soundtrack composition, we choose a temporal representation of video using sections, separated by sparse temporal boundaries of scene cuts (Cohen 1990; Prendergast 1992). To enable our music generator to anticipate future boundaries and align its output accordingly, we propose using boundary offsets. We conduct objective and subjective evaluations to measure the temporal synchronization between input video and generated music, and demonstrate that our approach outperforms dense temporal conditioning.
In addition, we tackle the challenge of using emotions as an intermediary to link music generation to video. Emotions are commonly represented with either categorical (Ekman 1971) or dimensional (continuous-valued) (Russell 1980) models. Based on prior findings from psychology and emotion research (Russell and Mehrabian 1977), we provide a mapping between categorical and dimensional emotion representations. This mapping enables the combined use of different emotion representations across data modalities, such as categorical emotions for video classification and dimensional emotions for MIDI generation, as in our case. This approach ensures flexibility and supports the use of training data labeled with either type of emotion representation.
We compare our method to state-of-the-art video-based MIDI generators through objective and subjective evaluation on different datasets, where it outperforms the state-of-the-art methods. Our main contributions are as follows:
@@
  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective and a majority of the objective metrics for all datasets.
@@
  • Based on psychological research, we establish a mapping between discrete emotion categories and continuous valence-arousal values, an enabling emotional connection between the input video and the generated music.
@@ 2 Related work
While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances. The Foley Music model (Gan et al. 2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network (Kipf and Welling 2017) and a Transformer (Vaswani et al. 2017). Similarly, Koepke et al. (2020) and Su et al. (2020) use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards. Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, however, these datasets typically contain fewer than 1k samples (Zhao et al. 2018; Li et al. 2018; Koepke et al. 2020; Su, Liu, and Shlizerman 2020). The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music (Su, Liu, and Shlizerman 2021).
@@
Some studies explore the more general task of generating symbolic music for arbitrary videos. The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing (Di et al. 2021). CMT employs the Lakh Pianoroll Dataset (Dong et al. 2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration (Hsiao et al. 2021). While this representation reduces sequence length, it significantly increases input dimensionality. Consequently, CMT uses a low temporal resolution of four ticks per beat, translating to a time resolution of 125 milliseconds for a song at 120 beats per minute. Furthermore, note events align precisely with beat subdivisions, whereas human performance naturally deviates from beat subdivisions due to expressive timing (Fabian, Timmers, and Schubert 2014). In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8 ms resolution. This resolution is sufficient to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions (Oore et al. 2020). Additionally, unlike CMT, our model integrates high-level emotional conditioning alongside low-level temporal features.
@@
In a follow-up to CMT, Zhuo et al. introduced the Symbolic Music Videos dataset, containing video-MIDI pairs (2023). They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples (Hawthorne et al. 2018). Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features (Radford et al. 2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
@@
The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning, but differs in application (Kang, Poria, and Herremans 2024). The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key. Their encoder-decoder transformer takes low- and high-level video features, along with user-provided primer chord and key, to generate chord sequences. These sequences are then arpeggiated using fixed patterns to create the final MIDI output. However, this method’s reliance on a fixed time grid omits the subtle expressive timing found in human performances. Additionally, its use of fixed arpeggiation patterns limits musical diversity, and requiring user-defined chord and key inputs restricts accessibility for non-musicians.
@@ 3.1 Conditional music generator
We extend an existing emotion-based MIDI generator trained with valence-arousal labels (Sulun, Davies, and Viana 2022) assigned to the Lakh MIDI dataset (Raffel 2016). By introducing temporal boundary conditioning and emotion mapping, we align music generation with input videos at inference time. Figure 2 illustrates the music generator, with the emotion-conditioning mechanism shown in the upper part and the boundary-conditioning mechanism in the lower part.
@@
Training dataset and preprocessing
We train our music generator on the Lakh Pianoroll Dataset (LPD) (Dong et al. 2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset (Raffel 2016). We tokenize the pianorolls using an event-based symbolic music representation (Oore et al. 2020). Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end. These tokens also encode pitch and instrument information. For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60. Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings (Dong et al. 2018). However, our method is adaptable to datasets with different instrument groupings.
TIMESHIFT tokens are used to move along the time axis, representing both note durations and the silences between them. Each token specifies a time increment in milliseconds. For example, an 800 ms shift is encoded as TIMESHIFT_800. We use a temporal resolution of 8 ms with a maximum shift of 1000 ms. Longer durations are represented using multiple consecutive TIMESHIFT tokens. We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches. We introduce the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions (Bharucha and Krumhansl 1983). However, this is a design choice, and our method can accommodate any selection of temporal boundaries.
To fully exploit available training data, we implement several strategies. Since we aim to generate multi-instrument compositions, we prioritize pieces with three or more instruments. However, one third of the Lakh MIDI dataset contains songs with only one or two instruments, with classical piano pieces dominating the single-instrument subset (Raffel 2016). As these samples still contain valuable information on musical structure, we avoid filtering them out. Instead, we prepend special tokens: FEWER_INSTRUMENTS for songs with two or fewer instruments and MORE_INSTRUMENTS for those with three or more. These tokens allow users to specify instrumentation preferences at inference time while leveraging the entire dataset during training. Additionally, to utilize MIDI samples without valence-arousal labels, we allow the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation. If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
@@
Temporal boundary matching
Our model generates music that is synchronized with the video’s temporal boundaries. We define temporal boundaries as scene cut locations in video and long-duration chord locations in music. Since we employ a hybrid approach, we first train a music generator capable of incorporating boundary locations as input and producing music with chords near these boundaries. To achieve this, we label the chords in the Lakh Pianoroll Dataset. During training, we consider only guitar and piano chords with at least three simultaneous notes that last for a minimum of two beats. These chords are labeled by inserting a CHORD token before the first ON token of each chord. To allow the model to generate chords not only at video scene cut locations but also independently where musically appropriate, we randomly remove 20% of CHORD tokens during training. An example of a labeled chord is shown in the top row of Figure 3. In training, the temporal locations of chords from the ground-truth MIDI serve as input boundaries, whereas during video-based inference, we replace these boundaries with video scene cut locations.
Chords are integral to a melody, providing harmonic and rhythmic support to surrounding notes, both preceding and following (Bharucha and Krumhansl 1983). During inference, forcing a CHORD token into the sequence at a specific location may cause the chord to sound off-beat or overly abrupt. This occurs because the model, having no prior knowledge of the upcoming chord, may generate preceding notes that do not align naturally with it. To address this, we develop a method that enables the model to “anticipate” upcoming chords and generate preceding notes and time shifts accordingly. Additionally, we train the model to generate the CHORD token itself, ensuring rhythmic consistency in the generated music. To achieve this, we define boundary offsets for each input token, representing the remaining time until the next boundary. These offsets are capped at a maximum value and, since our music generator is autoregressive, they are computed based on future chords rather than past ones. Furthermore, we amplify the loss of the CHORD token by a factor of 10, as it influences multiple preceding chord notes and plays a critical role in structuring the generated music. Sample offsets are illustrated in the bottom row of Figure 3.
We process boundary offsets through a feed-forward network, generating boundary offset encodings, which are then concatenated with learned positional encodings (Wang et al. 2020) along the feature dimension, as shown in the lower part of Figure 2. This architectural choice is motivated by several key factors. First, similar to the emotion-conditioning mechanism, we inject boundary offset encodings at the input level rather than within the transformer body, ensuring that the core model remains unchanged. This allows the model to process inputs with or without boundary offsets, enabling seamless fine-tuning by repeating the learned positional encodings along the feature dimension when boundary offsets are absent. Second, we avoid adding boundary offset encodings directly to learned positional encodings to maintain a distinction between the two. Finally, recent studies suggest that decoder-only transformers can implicitly learn positional encodings through their internal weights, even without explicitly adding them (Haviv et al. 2022; Kazemnejad et al. 2024). Based on this insight, we halve the feature length of learned positional encodings and allocate the remaining feature space to boundary offset encodings. The resulting vector sequence consists of positional encodings augmented with boundary offset encodings. Following the standard transformer model, we add this sequence to the token embeddings (Vaswani et al. 2017) before passing it to the transformer body with relative global attention (Huang et al. 2018). The transformer’s input is represented as:
where $x_{v}$ and $x_{a}$ are the valence and arousal inputs, $\mathbf{x}_{t}$ is the input token sequence, $\mathbf{b}$ represents the input boundary offsets, and $\mathbf{W_{pe}}$ is the learned positional encoding. The feed-forward networks $\text{FFN}_{v}$, $\text{FFN}_{a}$, and $\text{FFN}_{b}$ process the valence, arousal, and boundary offset inputs, respectively. The operations $\text{Concat}_{s}$ and $\text{Concat}_{f}$ denote concatenation along the sequence and feature dimensions, respectively.
@@
Algorithm 1 outlines the process of computing boundary offsets during inference, i.e., music generation. The model generates music autoregressively, producing one token per forward pass. We maintain a time cursor by tracking the generated TIMESHIFT tokens and compute the boundary offset, i.e., the time remaining until the next boundary, for each generated token. If the model generates a CHORD token, we calculate the absolute difference between the time cursor and each input boundary. The boundary with a difference smaller than the predefined sensitivity threshold of 1 s is considered successfully generated. We then remove these boundaries from future offset calculations by replacing them with infinity. Next, we compute the boundary offset for the generated token, regardless of its type. This offset is determined as the distance to the next closest boundary, capped by a maximum offset value. For simplicity, Algorithm 1 is presented for a single sample, but in practice, this operation is performed in minibatches. During training, we preprocess the entire input sequence at once by constructing a time grid instead of a time cursor, allowing us to calculate boundary offsets for all tokens simultaneously. For clarity, the initial list of generated tokens is shown as empty; however, in practice, we begin with a START token.
@@
[…]